i have a word problem that i need assisstance with, i am in an algebra level course.the word problem is below.
Two mechanics worked on a car. The first mechanic worked for 20 hours, and the second mechanic worked for 15 hours. Together they charged a total of $3950 . What was the rate charged per hour by each mechanic if the sum of the two rates was 225 per hour?

Answer by stanbon(75887)

Two mechanics worked on a car. The first mechanic worked for 20 hours, and the second mechanic worked for 15 hours. Together they charged a total of $3950 . What was the rate charged per hour by each mechanic if the sum of the two rates was 225 per hour?
-----
Equations:
r1 + r2 = 225
20r1 + 15r2 = 3950
---------------------------
Multiply thru 1st equation by 20 to get:
20r1 + 20r2 = 20*225
20r1 + 15r2 = 3950
========================
Subtract and solve for r2:
5r2 = 550
r2 = 110
---
Solve for "r1":
r1 + r2 = 225
r1 + 110 = 225
r1 = 115
===============
Cheers,
Stan H.